OpenStudy (lilsis76):

the sum of three consecutive terms in an arithmetic sequence is 30, and their product is 360. Find the three terms. Let x denote the middle term and d the common difference. well i got so far: ______+________+__________= 30 but ______x________x__________=360 ya, thats as far as I got and i am still stumped. I tried all kinds of things but i cant seem to get the right sum and multiplied values

OpenStudy (anonymous):

middle term must be 10

OpenStudy (lilsis76):

do i divide by 36? how do i know what to put in for the middle parts? i remember class mates saying something by going up or down by a k, or a 1 or something. is there a rule or something like that I have to know?

OpenStudy (anonymous):

ok, using your notation, x is the middle term, d is the common difference: (x-d) + x + (x+d) = 30. so what's x?

OpenStudy (lilsis76):

okay so the the left would be negative and the right would be positive.

OpenStudy (lilsis76):

okay @cnknd would it be 10-5 + 10 + 10+5 = 30?

OpenStudy (anonymous):

ugh... what???? ok take a step back, forget i told u that middle term is 10. look at the equation for the sum: (x-d) + x + (x+d) = 30 can u find x and d from this?

OpenStudy (lilsis76):

okay well i was thinking the 30/3 = the 10

OpenStudy (lilsis76):

@cnknd okay the d is the difference and x is the actual number. i am lost as to where to start

OpenStudy (anonymous):

(x-d) + x + (x+d) = 30 can you solve for either x or d?

OpenStudy (lilsis76):

i think we can solve the x. the 3x= 30 so x=10, right?

OpenStudy (anonymous):

yes but you can't really determine what d is though.

OpenStudy (anonymous):

but at least now you know why the x (the middle number) is 10.

OpenStudy (anonymous):

you can probably guess and check for the other 2 numbers (that's what i did)

OpenStudy (lilsis76):

okay, let me try and ill get back to you on this in a little

OpenStudy (lilsis76):

GOT IT! (10-8) +/* 10 +/* (10+8) +=30 * =360

OpenStudy (lilsis76):

THANK YOU SO MUCH @cnknd